### v1

# Continuous Optimization for Feature Selection with Permutation-Invariant Embedding and Policy-Guided Search

Abstract.
Feature selection removes redundant features to enhance both performance and computational efficiency in downstream tasks. Existing methods often struggle to capture complex feature interactions and adapt to diverse scenarios. Recent advances in this domain have incorporated generative intelligence to address these drawbacks by uncovering intricate relationships between features. However, two key limitations remain: 1) embedding feature subsets in a continuous space is challenging due to permutation sensitivity, as changes in feature order can introduce biases and weaken the embedding learning process; 2) gradient-based search in the embedding space assumes convexity, which is rarely guaranteed, leading to reduced search effectiveness and suboptimal subsets. To address these limitations, we propose a new framework that can: 1) preserve feature subset knowledge in a continuous embedding space while ensuring permutation invariance; 2) effectively explore the embedding space without relying on strong convex assumptions. For the first objective, we develop an encoder-decoder paradigm to preserve feature selection knowledge into a continuous embedding space. This paradigm captures feature interactions through pairwise relationships within the subset, removing the influence of feature order on the embedding. Moreover, an inducing point mechanism is introduced to accelerate pairwise relationship computations. For the second objective, we employ a policy-based reinforcement learning (RL) approach to guide the exploration of the embedding space. The RL agent effectively navigates the space by balancing multiple objectives. By prioritizing high-potential regions adaptively and eliminating the reliance on convexity assumptions, this search strategy effectively reduces the risk of converging to local optima. Finally, we conduct extensive experiments to demonstrate the effectiveness, efficiency, robustness and explicitness of our model. Our code and dataset are publicly accessible on GitHub111 https://anonymous.4open.science/r/CAPS-F01B.
isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Feature selection eliminates redundant and irrelevant features to improve both predictive performance and computational efficiency in downstream tasks. Despite the growing dominance of deep learning, feature selection remains indispensable in scenarios characterized by high-dimensional data, the need for interpretability, and limited resource constraints. For instance, in healthcare, identifying key biomarkers from extensive genetic datasets enhances model transparency and supports informed decision-making. In finance, extracting relevant features from transactional data facilitates the development of efficient and interpretable risk models. These examples demonstrate the importance of feature selection in addressing real-world challenges, even in the era of deep learning.
Existing works can be divided into three categories: 1) filter methods (Yang and Pedersen, 1997; Forman et al., 2003; Hall, 1999; Yu and Liu, 2003) rank features based on specific scoring criteria. They often select the top-K ranking features as the optimal subset; 2) wrapper methods (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997) iteratively generate and evaluate feature subsets. They select the feature subset that performs best on a downstream task as the optimal one; 3) embedded methods (Tibshirani, 1996; Sugumaran et al., 2007) integrate a feature selection regularization term (e.g., $L_{1}$ regularization) into model training. As the model converges, irrelevant features are penalized, and the non-zero coefficients indicate the optimal subset. While prior work has achieved notable success, it falls short in capturing complex feature interactions and adapting feature selection knowledge to diverse and dynamic scenarios.
Recently, the success of generative AI has inspired researchers to revisit the feature selection problem through a generative lens. By embedding discrete feature selection knowledge into a continuous embedding space, they not only capture complex feature interactions for generating enhanced feature subsets but also make the knowledge more adaptable to diverse scenarios (Xiao et al., 2023c; Ying et al., 2024a, 2023, b; Azim et al., 2024; Hu et al., 2024). However, there are two primary limitations:
  1. (1)
    Limitations 1: Permutation bias bring noise into the embedding space and lead to suboptimal performance. Existing methods fail to encode the fact that feature order does not impact model performance within the learned feature subset embedding. This oversight introduces bias into the embedding space, limiting the effectiveness of the search process in discovering optimal feature subsets.
  2. (2)
    Limitation 2: Convexity assumptions of the embedding space hinder effective exploration and search process. Prior literature often assumes that the embedding space is convex, with the expectation that gradient-based search will locate optimal solutions. In practice, this assumption rarely holds, which causes the search process to converge to suboptimal points and produce inferior feature subsets.
To address these limitations, we propose CAPS, a new framework that performs Continuous optimization for feAture selection by integrating Permutation-invariant embeddings with a policy-guided Search strategy. More specifically, given a large volume of feature selection records, where each record contains the indices of features within a subset and the corresponding model performance, we first design an encoder-decoder framework to capture the underlying patterns in the feature subset indices. To achieve permutation-invariant embeddings for feature subsets, we leverage a self-attention mechanism that symmetrically computes attention scores across all input feature indices. This design guarantees that any permutation of the input features produces identical embeddings, ensuring that the learned representations of feature subset indices are both robust and reliable. Considering the high computational complexity $O(N^{2})$ of pairwise attention calculation, we introduce a set of inducing points to alleviate the computational burden and accelerate processing. These inducing points act as intermediate representations, facilitating more efficient attention computations by reducing the need for full pairwise attention. This approach achieves a complexity of $O(NM)$, where $M$ denotes the number of inducing points, which is substantially smaller than $N$. After the encoder-decoder model converges, we deploy a policy-based reinforcement learning (RL) agent to explore the learned embedding space and identify the optimal feature subset. Based on model performance, we first select the top-K feature subsets as search seeds and input them into the well-trained encoder to obtain their corresponding embeddings. Next, we employ an RL agent to learn how to optimize these embeddings by maximizing the downstream task performance and minimizing the length of the feature subset. Throughout this process, the exploratory nature of the RL agent helps overcome the challenges of the non-convex embedding space, enabling the search for improved embeddings.
To summarize, our contributions are three-fold:
  • Problem: We investigate a new framework for automated feature selection from the generative lens. Remarkably, we propose to solve this problem by integrating permutation-invariant embedding and policy-guided search.
  • Algorithm: We design a novel encoder-decoder architecture to embed feature selection knowledge into a permutation-invariant space, removing permutation bias. We also employ a policy-based RL agent to explore the space for optimal solutions, overcoming strong convexity assumptions.
  • Evaluation: We conduct extensive experiments on 14 real-world datasets to validate the effectiveness of our approach. The experimental results show the superior performance of CAPS over the state-of-the-art feature selection methods.
2. Problem Statement
We propose an automated feature selection framework from a generative perspective, integrating permutation-invariant embeddings with a policy-guided search strategy to enhance effectiveness. Formally, we consider a dataset $\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$, where $\mathbf{X}$ denotes features, and $\mathbf{y}$ represents the corresponding labels. Using existing feature selection methods (Liu et al., 2019), we can gather $p$ feature selection records, denoted by ${\{(\mathbf{f}_{i},v_{i})\}}^{p}_{i=1}$, where each record consists of feature indices within the selected subset $\mathbf{f}_{i}$ and their corresponding model performance $v_{i}$. Then, we preserve feature selection knowledge into an embedding space $\mathcal{E}$ using a permutation-invariant encoder $\omega$ and decoder $\psi$, optimized by minimizing the reconstruction loss. Next, a policy-based reinforcement learning (RL) agent is utilized to explore the embedding space $\mathcal{E}$, aiming to identify the optimal embedding $\mathbf{E}^{*}$. This embedding can be used to reconstruct the optimal feature subset $\mathbf{f}^{*}$ by the decoder $\psi$, which maximizes downstream task performance $\mathcal{M}$. The optimization goal can be formulated as:
(1) $\mathbf{f}^{*}=\psi(\mathbf{E}^{*})=\mathrm{argmax}_{\mathbf{E}\in\mathcal{E}}\mathcal{M}(X[\psi(\mathbf{E})]),$
Finally, $\mathbf{f}^{*}$ is used to select the corresponding columns of $\mathcal{D}$, forming the optimal feature space for the downstream task.
3. Methodology
3.1. Framework Overview
Figure 1. An overview of CAPS. First, we develop an encoder-decoder structure to learn a permutation-invariant feature subset embedding space by optimizing the reconstruction loss (a). Then, we explore the learned embedding space through policy-guided RL search, aiming to maximize downstream task performance and minimize feature subset length (b). (c) and (d) illustrate the architecture of induced set attention block (ISAB) and pooling by multihead attention (PMA) respectively.
Figure 1 illustrates the overview of CAPS, including two main components: 1) permutation-invariant feature subset embedding (i.e., Figure 1 (a)); 2) policy-guided multi-objective search (i.e., Figure 1 (b)). Specifically, given a large collection of feature selection records, each record comprises a selected feature subset and the corresponding model performance. First, we train the permutation-invariant feature subset embedding module using these records. The encoder maps each feature subset to a continuous embedding by modeling pairwise relationships among feature IDs to ensure permutation invariance. The decoder then reconstructs the subset from the embedding to capture key feature patterns and interactions. After the learning converges, we select the top-K selection records from the collected data and regard them as search seeds. Then, a policy-based RL agent explores the learned continuous embedding space. The updated embedding is fed into the well-trained decoder to reconstruct the feature subset. The search process is guided by rewards that maximize downstream task performance and minimize the feature subset length. Finally, the feature subset achieving the highest performance is output as the optimal one.
3.2. Permutation-Invariant Embedding Learning
Why permutation-invariant embedding matters. Feature subsets are inherently permutation-invariant, as changes in the order of selected features do not impact the downstream task performance. Existing works (i.e., (Xiao et al., 2023c)) often use data augmentation strategy to alleviate this burden. In detail, they randomly shuffle the feature subset to increase the volume and diversity of the data, enhancing the embedding space. However, with large feature subsets, this strategy may leave many cases uncovered by random shuffling. To address this, we propose a permutation-invariant encoder-decoder.
Training data collection. To learn a feature subset embedding space, we collect $p$ feature selection records, denoted by $\{(\mathbf{f}_{i},v_{i})\}_{i=1}^{p}$, where $\mathbf{f}_{i}$ represents the feature indices of the $i$-th selected feature subset, and $v_{i}$ denotes the corresponding model performance. To achieve this, we use RL-based methods, specifically MARLFS (Liu et al., 2019), to automatically explore the feature space. In this method, each feature is controlled by an RL agent that decides to select or deselect it. The entire procedure is optimized to maximize downstream task performance. For more details, please refer to the referenced paper.
Leveraging a Permutation-Invariant Encoder-Decoder Framework to Embed Feature Subsets. In the following contents, we use the feature indices of a subset $\mathbf{f}$ as an example to illustrate the calculation process within the encoder-decoder structure.
Encoder $\omega$ : The encoder aims to learn a permutation-invariant mapping function $\omega$ that converts feature indices $\mathbf{f}\in\mathbb{R}^{1\times N}$ into continuous embedding $\mathbf{E}$, denoted by $\omega(\mathbf{f})=\mathbf{E}=[\mathbf{h_{1}},\mathbf{h_{2}},...,\mathbf{h_{N}}]\in\mathbb{R}^{N\times d}$, where $\mathbf{h_{n}}\in\mathbb{R}^{1\times d}$ is the $t$-th index embedding, $N$ is the number of input features, and $d$ is the hidden size of the embedding. Inspired by (Lee et al., 2019), we employ Multihead Attention Block (MAB) to encode pairwise interactions between features in the subset. The MAB is a modified version of the Transformer encoder block (Vaswani et al., 2017), excluding positional encoding and dropout. Formally, the MAB with parameters $\mathbf{W}$ is defined as:
(2) $MAB(\mathbf{Q},\mathbf{K},\mathbf{V})=LayerNorm(\mathbf{H}+rFF(\mathbf{H})),$
(3) $\mathbf{H}=LayerNorm(\mathbf{Q}+Multihead(\mathbf{Q},\mathbf{K},\mathbf{V};\mathbf{W})),$
where $rFF$ denotes row-wise feedforward layer that processes each row of feature indices independently and identically, $\mathbf{H}$ is an intermediate representation of $\mathbf{Q}$, $LayerNorm$ is layer normalization (Lei Ba et al., 2016), and $Multihead$ denotes the multihead attention mechanism. Here, $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ denotes query, key and value respectively. In particular, we set $\mathbf{Q}=\mathbf{K}=\mathbf{V}=\mathbf{f}$ to preserve complete original information and extract pairwise relationship among all features in the input subset. This design focuses on feature relationships rather than their order, thereby enhancing the robustness of the learned embeddings. However, the quadratic computational complexity $O(N^{2})$ of pairwise attention becomes prohibitive when the feature number $N$ increases. To mitigate the computational cost, we employ a set of inducing points as intermediate representations, encoding global information about the input features. Specifically, in our case, each distinct feature subset and its any permuted version are mapped to a tightly clustered region in the learned embedding space. The inducing points are then trained as a small set of representative anchors in that space, effectively capturing distinct patterns from these clusters and enabling efficient global aggregation. This strategy declines the computational complexity to $O(NM)$, where $M$ is the number of inducing points, which is significantly smaller than $N$. Formally, given feature indices $\mathbf{f}\in\mathbb{R}^{1\times N}$, the ISAB with $M$ inducing points $\mathbf{I}$ is defined as follows:
(4) $ISAB_{M}(\mathbf{f})=MAB(\mathbf{f},\mathbf{H},\mathbf{H})\in\mathbb{R}^{N\times d},$
(5) $\mathbf{H}=MAB(\mathbf{I},\mathbf{f},\mathbf{f})\in\mathbb{R}^{M\times d},$
where the inducing points $\mathbf{I}\in\mathbb{R}^{M\times d}$ are parameterized and jointly learned with the rest of the network, and $d$ is the hidden embedding size. The ISAB comprises of two MAB layers. It is analogous to low-rank projection or autoencoder models, where the input is first compressed to low dimensional representations and then reconstructed to the original dimension. Specifically, we replace the query of the first MAB with the inducing points $\mathbf{I}$ so that they can attend to the entire input features, producing a low-dimensional representation $\mathbf{H}$ that encodes global information from all input features. Then, we replace the key and value of the second MAB with $\mathbf{H}$ to inject the global information back into each feature. Finally, to capture high-ordered feature interactions, we stack two ISAB layers to construct our encoder:
(6) $\omega(\mathbf{f})=ISAB_{M}(ISAB_{M}(\mathbf{f}))=\mathbf{E}$
Decoder $\psi$ : The decoder $\psi$ aims to reconstruct the feature indices $\mathbf{f}$ from the continuous embedding $\mathbf{E}$ in the learned space $\mathcal{E}$, denoted by $\psi(\mathbf{E})=\mathbf{f}$. To efficiently aggregate information and facilitate the reconstruction process, we integrate MAB with learnable seed vectors $\mathbf{S}\in\mathbb{R}^{K\times d}$ which is called Pooling by Multihead Attention (PMA). These seed vectors serve as prototype queries, allowing each seed vector to selectively attend to embedding in the learned embedding space. In particular, each seed vector can be seen as pooling information from a different latent aspect of the embedding, thus enhancing the efficiency of information aggregation. Formally, given continuous embedding $\mathbf{E}\in\mathbb{R}^{N\times d}$, a PMA module with $K$ seed vectors is defined as follows:
(7) $PMA_{K}(\mathbf{E})=MAB(\mathbf{S},rFF(\mathbf{E}),rFF(\mathbf{E}))\in\mathbb{R}^{K\times d},$
where $\mathbf{E}$ denotes the input embedding to the decoder $\psi$, $rFF$ denotes row-wise feedforward layer, and $\mathbf{S}$ are parameterized and jointly learned with the rest of the network. Similar to the inducing points $\mathbf{I}$, we replace the query of the first MAB with the seed vectors $\mathbf{S}$ to ensure that they can pool information from different perspectives of $\mathbf{E}$. Then, to further capture the underlying interactions between $K$ outputs, we employ MAB and $rFF$ afterwards to construct our decoder:
(8) $\mathbf{\tilde{f}}=rFF(MAB(PMA_{K}(\mathbf{E})))\in\mathbb{R}^{1\times N},$
where $\mathbf{\tilde{f}}$ is the reconstructed feature indices, and it shares the same feature index dictionary with the input feature indices $\mathbf{f}$.
Optimization. To effectively reconstruct the feature indices $\mathbf{f}$, we optimize the encoder and decoder by minimizing the negative log-likelihood of $\mathbf{f}$ given $\mathbf{E}$, defined as follows:
(9) $\mathcal{L}_{rec}=-logP_{\psi}(\mathbf{f}|\mathbf{E})=-\sum^{N}_{n=1}logP_{\psi}(\mathbf{f}_{n}|\mathbf{h}_{n}).$
3.3. Policy-Guided Multi-Objective Search
Why selecting search seeds matters. As the model converges, we conduct a policy-guided RL search to identify the optimal feature subset embedding. Inspired by the initialization techniques in deep learning, we begin the search from advantageous points, referred to as search seeds, to accelerate and enhance performance. To obtain these seeds, we rank all collected records by model performance and select the top-K records to guide the search process.
Policy-guided multi-objective search. After constructing the embedding space, it must be explored to identify the optimal embedding for feature subset reconstruction. However, this space is not guaranteed to be convex. Moreover, the search for the optimal feature subset is inherently multi-objective, requiring trade-offs between maximizing downstream task performance and minimizing subset length. Thus, to conduct effective exploration and search process without strong convexity assumptions, we employ Proximal Policy Optimization (PPO) (Schulman et al., 2017) method. PPO is a widely used policy gradient multi-objective method due to its stable training process and strong empirical performance. It maintains policy stability by clipping and trust regions, effectively balancing exploration and exploitation in high-dimensional spaces without the reliance on convexity assumptions. Specifically, PPO aims to learn a policy $\pi$ to guide the exploration in the learned space by maximizing a clipped surrogate objective while constraining the step size of each policy update. This approach avoids excessively large updates and thereby mitigates the risk of collapse or divergence. To conduct effective exploration and search process, we elaborately design the components for the PPO RL agent. Given a feature subset embedding $\mathbf{E}\in\mathbb{R}^{N\times d}$, the PPO agent takes $\mathbf{E}$ as input, and outputs the enhanced embedding $\mathbf{E^{+}}$. Then, we feed $\mathbf{E^{+}}$ into the well-trained decoder $\psi$ to reconstruct the searched feature indices $\mathbf{f^{+}}$ as state $\mathbf{s}$ for the PPO agent. We use the downstream ML task performance of the searched feature indices $\mathbf{f^{+}}$ as reward $\mathbf{R}$. Specifically, the components in our RL framework include agent, state, and reward.
$Agent.$ In our design, the PPO agent $\mathcal{P}$ manipulates the embedding $\mathbf{E}$ to generate an enhanced embedding $\mathbf{E}^{+}$, expressed as: $\mathcal{P}(\mathbf{E})=\mathbf{E}^{+}$.
$State.$ The state $\mathbf{s}$ aims to describe the searched feature indices. We reconstruct the enhanced embedding $\mathbf{E}^{+}$ by utilizing the well-trained decoder $\psi$ to extract the representation of state, which can be denoted as:
(10) $\mathbf{s}=Rep(X[\mathbf{f}^{+}])=Rep(X[\psi(\mathbf{E}^{+})])=Rep(X[\psi(\mathcal{P}(\mathbf{E}))]),$
where $Rep$ is the one-hot encoding approach, $X$ is the given dataset.
Reward. To quantify the reward $\mathbf{R}$ obtained by the searched feature subset on the downstream ML task, we define a measurement as follows:
(11) $\mathbf{R}=\lambda(\mathcal{M}(X[\mathbf{f}^{+}])-\mathcal{M}(X[\mathbf{f}]))+(1-\lambda)\mathcal{N}[\mathbf{f}^{+}]),$
where $\mathcal{M}$ is the downstream ML task, $\mathbf{f}$ is the original selected feature indices, $\mathbf{f}^{+}$ is the searched feature indices, $\mathcal{N}[\mathbf{f}^{+}]$ denotes the length of $\mathbf{f}^{+}$, and $\lambda$ is the trade-off hyperparameter to balance the performance and the total length of the selected feature subset.
Solving the Optimization Problem. We train the PPO agent with multiple objectives. Specifically, we maximize a clipped surrogate objective, which is the discounted and cumulative reward during the iterative search process, while constraining the step size of each policy update. We encourage the agent to search more informative embeddings with higher model performance and fewer selected features. To accomplish this objective, we train the PPO agent in the actor-critic fashion. We define two separate networks for the actor and critic respectively and train them individually. We train the critic network by minimizing the difference between the cumulative discounted reward and their estimates, denoted as follows:
(12) $\mathcal{L}_{critic}=\frac{1}{T}\sum^{T}_{t=1}(V(\mathbf{s}_{t})-G_{t})^{2}=\frac{1}{T}\sum^{T}_{t=1}(V(\mathbf{s}_{t})-\sum^{T-t}_{k=0}\gamma^{k}R_{t+k})^{2},$
where $T$ is the total length of the trajectory, $\mathbf{s}_{t}$ is the state at $t$-th time step, $V(\mathbf{s}_{t})$ is the estimated output of the critic network, $G_{t}$ is the discounted cumulative reward at $t$-th time step, $\gamma\in[0\sim 1]$ is the discounted factor, and $R_{t}$ is the ground truth reward at $t$-th time step. To train the actor network, we maximize a clipped surrogate objective while constraining the step size of each policy update, denoted as follows:
(13) $\mathcal{L}_{actor}=\hat{\mathbb{E}}_{t}\left[min(r_{t}(\theta)\hat{A}_{t},clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})\right],$
where $\epsilon$ is a hyperparameter, $r_{t}(\theta)=\frac{\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})}{\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})}$ is the probability ratio of new policy $\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})$ and old policy $\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})$ at $t$-th time step, $\mathbf{a}_{t}$ is the action of the actor network at $t$-th time step, $\hat{A}_{t}=V(\mathbf{s}_{t})-G_{t}$ is the estimated advantage function at $t$-th time step, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)$ is the clipping function to remove the unexpected incentive for moving $r_{t}(\theta)$ outside of the interval [1-$\epsilon$, 1+$\epsilon$]. Eventually, by taking the minimum of the clipped and unclipped objective, we ensure that $r_{t}(\theta)$ is always in a trust region. Once the actor network converges, we obtain the optimal policy $\pi^{*}$ that chooses the most appropriate action for the exploration. In other words, it can guide the embedding $\mathbf{E}$ in the learned space $\mathcal{E}$ to shift toward the direction resulting in high downstream ML task performance without any convexity assumption. From the training settings of aforementioned PPO agent, the well-trained agent $\mathcal{P}$ guarantees that the enhanced embedding $\mathbf{E}^{+}$ contributes to higher downstream ML task performance. Therefore, to identify the optimal feature indices, we should reconstruct the feature subset indices $\mathbf{f}^{+}$ from the enhanced embedding $\mathbf{E}^{+}$. In particular, we feed the $K$ enhanced embedding $[\mathbf{E}^{+}_{1},\mathbf{E}^{+}_{2},...,\mathbf{E}^{+}_{K}]$ into the well-trained decoder $\psi$ to reconstruct the enhanced feature subset indices $[\mathbf{f}^{+}_{1},\mathbf{f}^{+}_{2},...,\mathbf{f}^{+}_{K}]$. Consequently, we utilize the reconstructed feature subset indices to generate different feature subsets from the original dataset. Then, we input them into the downstream ML task to evaluate their performance and choose the feature subset indices with the highest performance as the optimal one $\mathbf{f}^{*}$.
Table 1. Overall performance comparison. In this table, the best results are highlighted in $\bf bold$, and the second-best results are highlighted in $\underline{underlined}$. We evaluate classification (C), multi-classification (MC), and regression tasks in terms of F1-Score, Micro-F1, and 1-RAE respectively. (The higher the value is, the better the model performance is.)
Dataset Task #Samples #Features Original K-Best mRMR LASSO RFE LASSONet GFS SARLFS MARLFS RRA MCDM MEL GAINS CAPS
SpectF C 267 44 75.96 82.36 81.41 79.16 78.21 79.16 75.01 75.96 75.96 81.23 79.16 79.69 87.38 89.71
SVMGuide3 C 1243 21 77.81 77.02 77.42 74.77 78.90 76.51 83.70 78.92 82.26 77.63 77.26 73.95 83.68 84.22
German Credit C 1001 24 64.88 68.07 66.07 67.58 65.42 66.27 65.51 65.02 68.17 66.90 68.97 64.61 75.34 77.15
Credit Default C 30000 25 80.19 79.82 80.38 77.94 80.22 80.44 80.49 80.48 80.47 75.01 74.33 74.89 80.61 80.91
SpamBase C 4601 57 92.68 91.34 92.02 91.69 91.80 89.63 91.73 90.94 91.38 90.51 90.83 90.47 92.93 93.13
Megawatt1 C 253 38 81.60 80.08 80.08 83.78 80.08 82.83 78.60 82.75 78.71 86.05 82.83 78.18 90.42 91.83
Ionosphere C 351 34 92.85 92.74 89.93 94.14 95.69 94.14 92.77 89.82 88.51 92.72 92.74 90.16 97.10 98.55
Mice-Protein MC 1080 77 74.99 76.37 78.71 78.71 77.29 75.49 75.00 74.53 75.01 78.70 78.27 76.10 79.16 83.32
Coil-20 MC 1440 400 96.53 97.58 96.53 94.81 97.92 94.8 92.72 94.79 95.14 94.46 93.42 95.32 97.22 98.27
MNIST fashion MC 10000 784 80.15 79.45 80.40 79.50 80.70 79.45 80.15 80.10 79.90 79.95 79.75 79.98 81.00 81.15
UrbanSound MC 5702 16000 29.71 28.74 28.92 27.52 29.89 29.18 28.57 30.24 29.18 30.50 29.10 24.92 30.67 32.52
Openml_589 R 1000 25 50.95 55.30 55.28 59.74 55.03 56.32 59.74 53.43 53.51 54.54 55.43 53.99 59.76 59.77
Openml_616 R 500 50 15.63 24.86 24.27 29.52 23.90 22.98 44.44 25.35 25.68 24.67 24.28 42.26 47.39 48.52
IQ-Dataset R 20000 38 98.54 98.61 98.61 98.20 98.58 98.26 90.95 98.26 98.27 79.20 98.54 98.59 98.59 98.68
4. Experiments
(a) SVMGuide3
(b) German Credit
(c) Mice Protein
(d) Openml_616
Figure 2. The influence of data collection (CAPS-c), permutation invariance (CAPS-e) and RL search (CAPS-p) in CAPS.
4.1. Experimental Setup
4.1.1. Dataset Descriptions.
We employ 14 publicly accessible datasets from UCI (Public, 2022b), OpenML (Public, 2022a), CAVE (Nene et al., 1996), Kaggle (Howard, 2022), LibSVM (Chih-Jen, 2022), UrbanSound (Salamon et al., 2014) and OSF (Sajewicz-Radtke et al., 2024). These datasets can be categorized into 3 groups according to the types of ML tasks: 1) binary classification (C); 2) multi-class classification (MC); 3) regression (R). The statistics of these datasets can be found in Table 1.
4.1.2. Evaluation Metrics.
To conduct fair comparisons and alleviate the variance of the downstream ML model, we adopt Random Forest to assess the quality of selected feature subset and report the performance of each baseline algorithm by running five-fold cross-validation method. Moreover, we conduct all experiments via the hold-out setting to further obtain a fair comparison. For binary classification tasks, we use F1-score, Precision, Recall, and ROC/AUC. For multi-classification tasks, we use Micro-F1, Precision, Recall, and Macro-F1. For regression tasks, we use 1-Mean Average Error (1-MAE), 1-Mean Square Error (1-MSE), 1-Relative Absolute Error (1-RAE), and 1-Root Mean Square Error (1-RMSE). For all metrics, the higher the value is, the better the model performance is.
4.1.3. Baseline Algorithms.
We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point. (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning. Among the introduced baseline algorithms, K-Best and mRMR are categorized as filter methods. Lasso, RFE, and LassoNet are representative of embedded methods. GFS, SARLFS, and MARLFS are considered as wrapper methods. RRA, MCDM, and GAINS are classified as hybrid feature selection methods.
(a) SpectF
(b) German Credit
(c) SVMGuide3
(d) Openml_616
Figure 3. The visualization of original feature subset embeddings and permuted ones (from CAPS).
4.2. Experimental Results
4.2.1. Overall performance.
This experiment is conducted to assess the overall performance of our method. We compare the performance of our method and baseline algorithms. Table 1 shows the comparison of our method and 12 baseline algorithms in terms of F1-score, Micro-F1, and 1-RAE. The results demonstrate that our method outperforms other baseline algorithms across all domains and various tasks. There are two possible reasons for this observation: 1) The permutation-invariant encoder-decoder framework preserves feature subset knowledge in the learned embedding space, focusing on feature pairwise interactions rather than their order and removing the permutation bias in the embedding space; 2) The policy-based RL agent effectively explores the learned continuous embedding space and identifies the superior feature subset embeddings, overcoming the challenges of the non-convex space and converging to the global optimal feature subset embedding.
4.2.2. Ablation Study.
We conduct this experiment to study the impact of the data collection, permutation-invariant embedding, and policy-guided search on the model performance. We implement three model variants: 1) CAPS-c collects the feature subset-accuracy pairs randomly instead of employing basic and RL-based data collectors; 2) CAPS-e utilizes sequential encoder-decoder model to learn the continuous embedding space. 3) CAPS-p replaces the policy-guided search with Genetic Algorithm (GA) to explore the embedding space and identify the optimal feature subset embedding. We randomly choose four datasets to conduct this experiments, which are SVMGuide3, German Credit, Mice Protein, and Openml_616. For each type of task, we evaluate the model performance from four different perspectives. For binary classification tasks, we use Precision, Recall, F1-score, and ROC/AUC. For multi-classification tasks, we use Precision, Recall, Micro-F1, and Macro-F1. For regression tasks, we use 1-MAE, 1-MSE, 1-RAE, and 1-RMSE. Figure 2 shows the overall experiment results. We observe that the performance of CAPS is higher than CAPS-c, CAPS-e and CAPS-p on all datasets. The potential reasons for this observation are: 1) the feature selection records generated by the data collectors are more accurate and robust, constructing a more comprehensive and effective continuous embedding space for searching the optimal feature subset; 2) sequential encoder-decoder model fails to embed permutation invariance in the embedding space, introducing permutation bias and leading to local optima; 3) policy-guided RL agent conducts effective exploration for optimal feature subset, eliminating the reliance on convexity assumptions and avoiding convergence to local optima. Thus, this experiment demonstrates the significance of data collector, permutation-invariant embedding and policy-guided search for enhancing the performance of CAPS.
(a) Precision
(b) Recall
(c) F1 Score
(d) ROC/AUC
Figure 4. Comparison of random search seeds, and top-K search seeds in terms of Precision, Recall, F1 Score, and ROC/AUC.
(a) RandomForest
(b) XGBoost
(c) SVM
(d) KNN
(e) Decision Tree
Figure 5. Comparison of different downstream ML models in terms of Micro-F1 on UrbanSound.
Figure 6. Comparison of the selected feature subset size of our method, the best baseline algorithm (Second Best), and the original feature set (Original).
(a) Original Feature Set
(b) CAPS-selected Feature Subset
Figure 7. Case Study. Comparison of traceability on the original feature set and selected feature subset.
4.2.3. Study of the permutation sensitivity of learned feature subset embeddings.
We conduct this experiment to study the permutation sensitivity of learned feature subset embeddings. We randomly choose four datasets and visualize the embeddings of the original feature subsets along with their corresponding permuted versions. In detail, we first randomly collect five distinct feature subset as the original representatives. Each feature subset is subjected to a permutation operation for 20 times to generate the corresponding permuted representatives. Then, we employ the well-trained encoder to obtain the embeddings for both the original and permuted feature subsets. Subsequently, we use T-SNE to map these embeddings into a 2-dimensional space for visualization. Figure 3 shows the visualization results, in which each color represents a unique feature subset along with its corresponding permuted versions. The embeddings of the original feature subset and the corresponding permuted versions are represented by pentagrams and squares respectively. We find that the distribution locations of the permuted feature subset embeddings are consistently clustered around their corresponding original feature subset embedding. A potential reason for this observation is that the learned continuous embedding space inherently captures permutation invariance, ensuring that variations in feature order do not impact the encoded embedding distribution. Therefore, this experiment demonstrates that our proposed encoder-decoder embeds feature subset knowledge into a permutation-invariant continuous embedding space, successfully removing the permutation bias.
4.2.4. Study of the impact of search seeds.
We conduct this experiment to study the impact of initial search seeds for the RL-based search process. To mitigate bias, we conduct the comparison on eight randomly chosen datasets and in term of four metrics, which are Precision, Recall, F1 Score, and ROC/AUC. We compare the search process using random K historical records with the one using top-K historical records based on model performance. Figure 4 shows the overall comparison results. The experimental results demonstrate that the search process with top-K records as initial search seeds consistently outperforms the one without across all metrics. A plausible explanation for this observation is that the starting of the search process is sensitive to the initial starting points. Initializing the search with top-K historical records not only provides a more promising starting region in the embedding space but also reduces the likelihood of exploring less relevant or suboptimal areas. The RL-based search agent can leverage the historical information to form a lower bound and focus on refining already competitive solutions. This leads to faster convergence and more stable performance across various datasets. Another noteworthy observation is that using top-K seeds helps to alleviate the risk of becoming trapped in local optima since these seeds often span multiple good regions in the embedding space. In contrast, random seeds might guide the search process toward less advantageous regions, increasing the variance in final performance. Therefore, these observations demonstrate that search seeds are crucial for the RL agent to explore the space and identify the optimal feature subset akin to the significance of initialization for deep neural networks.
4.2.5. Study of the robustness of CAPS over downstream ML tasks.
We conduct this experiment to assess the robustness of CAPS by changing the downstream ML model to Random Forest (RF), XGBoost (XGB), Support Vector Machine (SVM), K-Nearest Neighborhood (KNN), and Decision Tree (DT) respectively. Figure 5 shows the comparison results on UrbanSound dataset in terms of Micro-F1. The results demonstrate that CAPS outperforms other baselines regardless of downstream ML models. A potential reason for this observation is that the data collector can customize the feature selection records based on the downstream ML task. Then, the encoder-decoder model is capable of embedding the preference and properties of the ML model into the continuous embedding space. Finally, by leveraging these information, the RL agent can reach the global optimal embedding point instead of local optima. Thus, this experiment successfully demonstrates the robustness of CAPS.
4.2.6. Study of the size of selected feature subsets.
We conduct this experiment to assess the ability of the RL agent to optimize multiple objectives simultaneously, improving model performance while reducing selected features. We conduct the comparison on six randomly selected datasets, and compare the size of selected feature subset of our method with the best-performing baseline algorithm and the original feature set. Figure 6 demonstrates the results of overall comparison in terms of the ratio of selected feature subset size to original feature set size. The ratio indicates the proportion of selected features relative to the original features. The results show that the selections of our method are consistently smaller than the ones selected by the second-best baseline algorithms on all the datasets. Although our selected feature subsets contain noticeably fewer features, they consistently achieve superior or comparable model performance. A potential reason for this observation is that the policy-guided RL agent is capable of optimizing the learned embeddings by maximizing the downstream task performance and minimizing the length of the selected subset. Therefore, this experiment demonstrates that the policy-guided RL agent excels at striking a balance between model performance and feature-efficiency.
4.2.7. Case study.
We conduct this experiment to examine the traceability of CAPS. We rank the top 7 significant features for prediction in both the original feature set and the selected feature subset using the IQ-Dataset. Figure 7 (a-b) show the top 7 significant features among 38 original features and the selected features of CAPS respectively. The predictive label is the summation of NVER_SC_119 and VERB_SC_119, which is the summed standardized score from 1 to 19 for the Nonverbal Intelligence scale and the summed standardized score from 1 to 19 for the Verbal Intelligence scale respectively. Figure 7 (a) illustrates that these two crucial features are absent from the top 7 features among the original 38 features. However, from Figure 7 (b), we observe that CAPS successfully selected these two features from 38 original features. A potential reason for this observation is that CAPS is capable of capturing the complicated interactions among features as well as the causality between features and predictive label. Therefore, this experiment demonstrates the traceability of the feature subset selected by CAPS, highlighting its effectiveness from a different perspective.
5. Related Works
Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time. Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005). However, such methods generally ignore the dependency and interactions among features. 2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996). However, such methods require excessive model training, resulting in high computational costs. 3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006). However, such methods heavily rely on the underlying model structure. Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a). For instance, (Xiao et al., 2023c) introduces a framework that encompasses a sequential encoder, an accuracy evaluator, a sequential decoder, and a gradient ascent optimizer, embedding discrete feature selection knowledge into a continuous embedding space and conducting gradient-ascent search process for the optimal feature subset. However, this approach is limited by: 1) incapable of encoding the permutation invariance into the embedding space; 2) the reliance on the convexity assumptions about the feature subset embedding space. To address these drawbacks, CAPS combines permutation-invariant embedding and policy-guided search. The permutation-invariant embedding reduces dependency on feature order, creating a more robust embedding space that preserves feature selection knowledge. The policy-guided search applies a reinforcement learning strategy to effectively explore this space, discovering better feature subsets without relying on strong convexity assumptions.
6. Conclusions
In this paper, we propose a generative automated feature selection framework that integrates permutation-invariant embeddings and policy-guided search to overcome permutation bias and convexity assumptions in existing methods. In detail, we first train a permutation-invariant feature subset embedding module using feature selection records by optimizing the reconstruction loss. The encoder maps feature subsets to continuous embeddings, while the decoder reconstructs the subsets from these embeddings. Moreover, considering the high computational cost of the pairwise interaction calculations, we leverage a set of inducing points as intermediate representations. These inducing points are expected to extract information about the input feature subset, contributing to capture global feature pattern and facilitate more efficient attention calculations by reducing the need for full pairwise attention. As the model converges, we employ a policy-guided RL agent to explore the learned space for superior feature subsets, overcoming the reliance on convex assumptions and mitigating the risk of converging to local optima. Finally, extensive experiments demonstrate that our framework offers significant insights: 1) Feature order sensitivity introduces permutation bias into the embedding space and undermines the embedding search process. 2) Policy-guided RL search prioritizes high-potential regions and eliminates the reliance on convexity assumptions, avoiding convergence to local optima. These insights emphasize the potential of reducing permutation sensitivity and overcoming non-convexity to improve the effectiveness and efficiency of automated feature selection. In the future, a promising direction will be to enhance the efficiency of the embedding search process. This can be achieved by reducing its strong reliance on a well-trained decoder for subset reconstruction.

### v2

@@
 Finally, we conduct extensive experiments to demonstrate the effectiveness, efficiency, robustness and explicitness of our model.
-Our code and dataset are publicly accessible on GitHub111 https://anonymous.4open.science/r/CAPS-F01B.
-isbn: 978-1-4503-XXXX-X/2018/06
+Our code and dataset are publicly accessible on GitHub [^1] https://github.com/RayLiu1103/CAPS.
+Automated Feature Selection; Representation Learning; Reinforcement Learning
+journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736891isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Feature selection
 1. Introduction
@@
 Reward. To quantify the reward $\mathbf{R}$ obtained by the searched feature subset on the downstream ML task, we define a measurement as follows:
-(11) $\mathbf{R}=\lambda(\mathcal{M}(X[\mathbf{f}^{+}])-\mathcal{M}(X[\mathbf{f}]))+(1-\lambda)\mathcal{N}[\mathbf{f}^{+}]),$
+(11) $\mathbf{R}=\lambda(\mathcal{M}(X[\mathbf{f}^{+}])-\mathcal{M}(X[\mathbf{f}]))+(1-\lambda)\mathcal{N}[\mathbf{f}^{+}],$
 where $\mathcal{M}$ is the downstream ML task, $\mathbf{f}$ is the original selected feature indices, $\mathbf{f}^{+}$ is the searched feature indices, $\mathcal{N}[\mathbf{f}^{+}]$ denotes the length of $\mathbf{f}^{+}$, and $\lambda$ is the trade-off hyperparameter to balance the performance and the total length of the selected feature subset.
@@
 (13) $\mathcal{L}_{actor}=\hat{\mathbb{E}}_{t}\left[min(r_{t}(\theta)\hat{A}_{t},clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})\right],$
-where $\epsilon$ is a hyperparameter, $r_{t}(\theta)=\frac{\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})}{\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})}$ is the probability ratio of new policy $\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})$ and old policy $\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})$ at $t$-th time step, $\mathbf{a}_{t}$ is the action of the actor network at $t$-th time step, $\hat{A}_{t}=V(\mathbf{s}_{t})-G_{t}$ is the estimated advantage function at $t$-th time step, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)$ is the clipping function to remove the unexpected incentive for moving $r_{t}(\theta)$ outside of the interval [1-$\epsilon$, 1+$\epsilon$].
+where $\epsilon$ is a hyperparameter, $r_{t}(\theta)=\frac{\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})}{\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})}$ is the probability ratio of new policy $\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})$ and old policy $\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})$ at $t$-th time step, $\mathbf{a}_{t}$ is the action of the actor network at $t$-th time step, $\hat{A}_{t}=V(\mathbf{s}_{t})-G_{t}$ is the estimated advantage function at $t$-th time step, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)$ is the clipping function to remove the unexpected incentive for moving $r_{t}(\theta)$ outside of the interval $\epsilon, 1$$\epsilon$].
 Eventually, by taking the minimum of the clipped and unclipped objective, we ensure that $r_{t}(\theta)$ is always in a trust region.
@@
 Therefore, to identify the optimal feature indices, we should reconstruct the feature subset indices $\mathbf{f}^{+}$ from the enhanced embedding $\mathbf{E}^{+}$.
-In particular, we feed the $K$ enhanced embedding $[\mathbf{E}^{+}_{1},\mathbf{E}^{+}_{2},...,\mathbf{E}^{+}_{K}]$ into the well-trained decoder $\psi$ to reconstruct the enhanced feature subset indices $[\mathbf{f}^{+}_{1},\mathbf{f}^{+}_{2},...,\mathbf{f}^{+}_{K}]$.
+In particular, we feed the $K$ enhanced embedding $\mathbf{E}^{+}_{1},\mathbf{E}^{+}_{2},...,\mathbf{E}^{+}_{K}$] into the well-trained decoder $\psi$ to reconstruct the enhanced feature subset indices $\mathbf{f}^{+}_{1},\mathbf{f}^{+}_{2},...,\mathbf{f}^{+}_{K}$].
 Consequently, we utilize the reconstructed feature subset indices to generate different feature subsets from the original dataset.
@@
 4.1.3. Baseline Algorithms.
-We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point.
+We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021a) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point.
 (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning.
@@
 Figure 3. The visualization of original feature subset embeddings and permuted ones (from CAPS).
+4.1.4. Experimental Settings.
+We conducted all experiments on the Windows 11 operating system, AMD Ryzen 5 5600X CPU, and NVIDIA GeForce RTX 3070Ti GPU, with the framework of Python 3.10.15 and PyTorch 2.5.1 (Paszke et al., 2019).
+4.1.5. Hyperparameters and Reproducibility.
+To collect potential feature subsets and their corresponding model performance, we run MARLFS for 300 epochs.
+To increase the diversity of training data, we augment the collected feature subset-accuracy records by permuting the order of each feature subset indices 25 times.
+The Encoder comprises of two layers of ISAB.
+The Decoder comprises of 3 components, which are PMA, MAB and rFF.
+The number of multi-attention head is 4.
+The embedding size of each feature index is 128.
+To train the encoder and decoder, we set batch size as 64, the step size as 0.001, and the dimension of seed vectors and inducing points as 32 respectively.
+During the search process, we used the top 25 feature selection records as the initial search seeds to search for the optimal embedding vector.
+To conduct stable and robust search, we set search epoch as 10, batch size as 512, learning rate of the actor as 0.0003, learning rate of the critic as 0.001, reward trade-off as 0.1, reward discounted factor as 0.99, search step as 1000, and the clipping ratio as 0.2 respectively.
+(a) Precision
+(b) Recall
+(c) F1 Score
+(d) ROC/AUC
+Figure 4. Comparison of random search seeds, and top-K search seeds in terms of Precision, Recall, F1 Score, and ROC/AUC.
+(a) RandomForest
+(b) XGBoost
+(c) SVM
+(d) KNN
+(e) Decision Tree
+Figure 5. Comparison of different downstream ML models in terms of Micro-F1 on UrbanSound.
 4.2. Experimental Results
@@
 Thus, this experiment demonstrates the significance of data collector, permutation-invariant embedding and policy-guided search for enhancing the performance of CAPS.
-(a) Precision
-(b) Recall
-(c) F1 Score
-(d) ROC/AUC
-Figure 4. Comparison of random search seeds, and top-K search seeds in terms of Precision, Recall, F1 Score, and ROC/AUC.
-(a) RandomForest
-(b) XGBoost
-(c) SVM
-(d) KNN
-(e) Decision Tree
-Figure 5. Comparison of different downstream ML models in terms of Micro-F1 on UrbanSound.
-Figure 6. Comparison of the selected feature subset size of our method, the best baseline algorithm (Second Best), and the original feature set (Original).
-(a) Original Feature Set
-(b) CAPS-selected Feature Subset
-Figure 7. Case Study. Comparison of traceability on the original feature set and selected feature subset.
 4.2.3. Study of the permutation sensitivity of learned feature subset embeddings.
@@
 Thus, this experiment successfully demonstrates the robustness of CAPS.
+Figure 6. Comparison of the selected feature subset size of our method, the best baseline algorithm (Second Best), and the original feature set (Original).
+(a) Clipping Trade-off
+(b) Reward Trade-off
+Figure 7. The hyperparameter sensitivity test on SpectF.
 4.2.6. Study of the size of selected feature subsets.
@@
 Therefore, this experiment demonstrates that the policy-guided RL agent excels at striking a balance between model performance and feature-efficiency.
+4.2.7. Study of the Hyperparameter Sensitivity of Search Process.
+We conduct this experiment to study the sensitivity of the search process performance to the the clipping trade-off $\epsilon$ and the reward trade-off $\lambda$.
+As part of the surrogate objective, the clipping trade-off $\epsilon$ can stabilize training and prevent excessively large updates to the original policy.
+Small value of $\epsilon$ results in stricter policy updates, ensuring that each step is small and stable, but at the risk of limiting exploration and making it harder for the policy to improve.
+Conversely, larger value of $\epsilon$ results in more generative policy updates, facilitating quicker exploration of the policy space.
+In other words, a smaller $\epsilon$ prioritizes exploitation over exploration, and vice versa.
+The reward trade-off $\lambda$ is designed for the RL agent to balance the attention on the feature subset performance and the total length of the subset from equation 11.
+In particular, a larger $\lambda$ prioritizes the feature subset performance over total length, and vice versa.
+We set both $\epsilon$ and $\lambda$ vary from 0.1 to 0.9, then train the CAPS on Spectf.
+The model performance is shown in Figure 7 in terms of Precision, Recall and F1-Score.
+Overall, we observed that CAPS achieves best performance when $\epsilon$ is set to 0.2 and $\lambda$ is set to 0.1.
+From an empirical perspective, PPO has been shown to perform robustly across a wide range of tasks, with $\epsilon$ typically set to 0.2 in many standard tasks (Schulman et al., 2017), which aligns with our findings.
+As for the reward trade-off $\lambda$, the model performance remains relatively stable across most values of $\lambda$ and occasionally increases when $\lambda$ is set to 0.1.
+These findings provide insights into how $\epsilon$ and $\lambda$ influence the search results and how we should choose optimal values for them.
 4.2.8. Case study.
@@
 Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time.
-Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005).
+Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005; Biesiada and Duch, 2008; Ding et al., 2014).
 However, such methods generally ignore the dependency and interactions among features.
-2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996).
+2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Liu et al., 2021b; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996).
 However, such methods require excessive model training, resulting in high computational costs.
-3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006).
+3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006; Kumagai et al., 2022; Koyama et al., 2022).
 However, such methods heavily rely on the underlying model structure.
@@
 However, this approach is limited by: 1) incapable of encoding the permutation invariance into the embedding space; 2) the reliance on the convexity assumptions about the feature subset embedding space.
-To address these drawbacks, CAPS combines permutation-invariant embedding and policy-guided search.
-The permutation-invariant embedding reduces dependency on feature order, creating a more robust embedding space that preserves feature selection knowledge.
-The policy-guided search applies a reinforcement learning strategy to effectively explore this space, discovering better feature subsets without relying on strong convexity assumptions.
+To address these drawbacks, we integrate a permutation-invariant embedding mechanism with a policy-guided search strategy in CAPS.
+The permutation-invariant embedding is designed to mitigate the sensitivity to feature ordering, constructing a more robust embedding space that preserves feature selection knowledge.
+The policy-guided search leverages a reinforcement learning strategy to effectively explore this space, identifying better feature subsets without relying on strong convexity assumptions.
+(a) Original Feature Set
+(b) CAPS-selected Feature Subset
+Figure 8. Case Study. Comparison of traceability on the original feature set and selected feature subset.
 6. Conclusions

### v3

@@
-Continuous Optimization for Feature Selection with Permutation-Invariant Embedding and Policy-Guided Search
+arXiv:2505.11601v3[cs.LG] 27 Feb 2026
 Abstract.
@@
 These examples demonstrate the importance of feature selection in addressing real-world challenges, even in the era of deep learning.
-Existing works can be divided into three categories: 1) filter methods (Yang and Pedersen, 1997; Forman et al., 2003; Hall, 1999; Yu and Liu, 2003) rank features based on specific scoring criteria.
+Existing works can be divided into three categories: 1) filter methods (Yang and Pedersen, 1997; Forman and others, 2003; Hall, 1999; Yu and Liu, 2003) rank features based on specific scoring criteria.
 They often select the top-K ranking features as the optimal subset; 2) wrapper methods (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997) iteratively generate and evaluate feature subsets.
@@
 4.1.3. Baseline Algorithms.
-We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021a) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point.
+We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021b) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point.
 (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning.
@@
 4.1.4. Experimental Settings.
-We conducted all experiments on the Windows 11 operating system, AMD Ryzen 5 5600X CPU, and NVIDIA GeForce RTX 3070Ti GPU, with the framework of Python 3.10.15 and PyTorch 2.5.1 (Paszke et al., 2019).
+We conducted all experiments on the Windows 11 operating system, AMD Ryzen 5 5600X CPU, and NVIDIA GeForce RTX 3070Ti GPU, with the framework of Python 3.10.15 and PyTorch 2.5.1.
 4.1.5. Hyperparameters and Reproducibility.
@@
 Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time.
-Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005; Biesiada and Duch, 2008; Ding et al., 2014).
+Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman and others, 2003; Yu and Liu, 2003; Peng et al., 2005; Biesiada and Duch, 2008; Ding et al., 2014).
 However, such methods generally ignore the dependency and interactions among features.
-2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Liu et al., 2021b; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996).
+2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Liu et al., 2021a; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996).
 However, such methods require excessive model training, resulting in high computational costs.
@@
 However, such methods heavily rely on the underlying model structure.
-Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a).
+Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a; Wang et al., 2026; Liu et al., 2025, 2026b, 2026a).
 For instance, (Xiao et al., 2023c) introduces a framework that encompasses a sequential encoder, an accuracy evaluator, a sequential decoder, and a gradient ascent optimizer, embedding discrete feature selection knowledge into a continuous embedding space and conducting gradient-ascent search process for the optimal feature subset.